Outline Objective Purpose

Background

Data

Regression Time series

ARIMA CIR

Interest Rate Modeling

Jaime Frade

April 18, 2008

Outline Objective Purpose Background

Data

Regression Time series

ARIMA CIR

Objective Purpose Background Data

3 month LIBOR 3 month T-rate 30 year T-rate

Regression Time series Model Forcasted value

ARIMA Model Forcasted value

CIR Model Forcasted value

Jaime Frade  Interest Rate Modeling

Outline Objective Purpose

Background

Data

Regression Time series

ARIMA CIR

Outline Objective Purpose

Background

Data

Regression Time series

ARIMA CIR

Outline Objective Purpose

Background

Data

Regression Time series

ARIMA CIR

Outline Objective Purpose

Background

Data

Regression Time series

ARIMA CIR

Outline Objective Purpose

Background

Data

Regression Time series

ARIMA CIR 30 year T-rate

Outline Objective Purpose

Background

Data

Regression Time series

ARIMA CIR Forcasted value

Outline Objective Purpose

Background

Data

Regression Time series

ARIMA CIR Forcasted value

Outline Objective Purpose

Background

Data

Regression Time series

ARIMA CIR Forcasted value